# Guaranteed Generation from Large Language Models

### v1

@@ A.4 Characterization of $g^{\prime}$
A.5 Relationship between $\text{KL}(g^{\prime}||a^{\prime})$ and Acceptance Rate $AR_{a^{\prime}}$
By the same reasoning as in App. A.4, applied to $a^{\prime}$, we see that $g^{\prime}(y)\propto a^{\prime}(y)b(y)$, and therefore $g^{\prime}(y)=\frac{1}{Z^{\prime}}a^{\prime}(y)b(y)$ where $Z^{\prime}=\sum_{y\in{\mathcal{Y}}}a^{\prime}(y)b(y)$.
@@
A.6 Core theorem about GUARD
@@
Appendix B GUARD’s training algorithm
@@
2:
3$\triangleright$ Initialization with CAP proposal
4:$\pi_{\theta}\leftarrow a(\cdot)$triangleright Policy
@@
Appendix C Additional metrics details
@@
In Theorem 2, we aim to minimize the KL divergence between the gold distribution $g$ and the distribution $g^{\prime}$ produced by our framework. We estimate the quality of GUARD, represented by $\text{KL}(g||g^{\prime})$, through the following method:
$\text{KL}(g||g^{\prime})=\mathbb{E}_{y\sim g}\log\frac{g(y)}{g^{\prime}(y)}$ $=\mathbb{E}_{y\sim g}\log\frac{a(y)b(y)}{Z}\frac{Z^{\prime}}{a^{\prime}(y)b(y)}$
$=\mathbb{E}_{y\sim g}\log\frac{a(y)}{a^{\prime}(y)}-\log\frac{Z}{Z^{\prime}}.$
We measured distances using the disco [18] toolkit. [^15] https://github.com/naver/disco
@@
Appendix D Lexical constraint experiment
@@
Appendix E Sentiment reversal experiment
@@
Appendix F Discussion of additional sampling techniques: IMH and QRS
@@ F.2 Experimental results
In the case of QRS and especially IMH, the quality gains come at a high cost in terms of additional inference time: to halve the projected KL divergence — which is a lower bound of the true KL divergence, see Footnote 21 — QRS requires 1.36 times the inference cost of Algorithm 1, while IMH requires 10 times the cost. Moreover, to reduce the KL divergence to 10% of the original KL, QRS uses almost 10 times the inference cost of Algorithm 1, while IMH consumes 50 times this cost. This experiment demonstrates that the random walk nature of MCMC methods can struggle to efficiently mimic the gold distribution. This result highlights why accurately approximating $g$ with the proposal distribution is crucial: it can lead to more efficient sampling that diverges less from the target distribution without incurring the extreme computational costs associated with methods like IMH.

### v2

@@ B.4 Characterization of $g^{\prime}$
B.5 Relationship between $\text{KL}(g^{\prime}||a^{\prime})$ and Acceptance Rate $AR_{a^{\prime}}$
By the same reasoning as in App. B.4, applied to $a^{\prime}$, we see that $g^{\prime}(y)\propto a^{\prime}(y)b(y)$, and therefore $g^{\prime}(y)=\frac{1}{Z^{\prime}}a^{\prime}(y)b(y)$ where $Z^{\prime}=\sum_{y\in{\mathcal{Y}}}a^{\prime}(y)b(y)$.
@@
B.6 Core theorem about GUARD
@@
B.7 Minimization of $\text{KL}(g||a^{\prime})$
As discussed in Section 3 and in the interpretation of Theorem 2, minimizing $\text{KL}(g||a^{\prime})$ results in finding a suitable $a^{\prime}$ for GUARD. In this section, we show that both the SFT and DPG approximation techniques seek to minimize this objective.
The SFT loss corresponds to the cross-entropy of the model $a^{\prime}$ to be learned, using samples from $g$, i.e., $-\mathbb{E}_{y\sim g}\log a^{\prime}(y)$. Minimizing this term with respect to $a\textquoteright$ ’ is equivalent to minimizing $\text{KL}(g||a^{\prime})$ since $\text{KL}(g||a^{\prime})=\mathbb{E}_{y\sim g}\log\frac{g(y)}{a^{\prime}(y)}=\mathbb{E}_{y\sim g}\log g(y)-\mathbb{E}_{y\sim g}\log a^{\prime}(y)=H_{g}-\mathbb{E}_{y\sim g}\log a^{\prime}(y)$ where $H_{g}$ is the entropy of $g$, which is a constant independent of $a\textquoteright$ ’.
The objective of DPG is also to minimize the KL divergence (or equivalently, the cross-entropy) between the target distribution $g$ and the learned policy $a^{\prime}=\pi_{\theta}$ as detailed in Parshakova et al. [37, Section 3.2]. The derivation can be summarized as follows: $\nabla_{\theta}\text{KL}(g||\pi_{\theta})=\nabla_{\theta}(H_{g}-\mathbb{E}_{y\sim g}\log\pi_{\theta}(y))=-\mathbb{E}_{y\sim g}\nabla_{\theta}\log\pi_{\theta}(y)=-\mathbb{E}_{y\sim\pi_{\theta}}\frac{g(y)}{\pi_{\theta}(y)}\nabla_{\theta}\log\pi_{\theta}(y)$, where the last step is obtained by applying importance sampling using $\pi_{\theta}$ as the proposal. Performing gradient descent on $-\mathbb{E}_{y\sim\pi_{\theta}}\frac{g(y)}{\pi_{\theta}(y)}\nabla_{\theta}\log\pi_{\theta}(y)$ leads to the DPG update rule shown at Line 21 of Algorithm 2 (see below).
@@
2:$\pi_{\theta}\leftarrow a(\cdot)$triangleright Policy
3:
4$\triangleright$ Initialization with CAP proposal
@@ Appendix C GUARD’s Training Algorithm
Appendix D Additional Metrics Details
@@
In Theorem 2, we aim to minimize the KL divergence between the gold distribution $g$ and the distribution $g^{\prime}$ produced by our framework. We can estimate the quality of GUARD, represented by $\text{KL}(g||g^{\prime})$, by using the following identity:
$\text{KL}(g||g^{\prime})$ $=\mathbb{E}_{y\sim g}\log\frac{g(y)}{g^{\prime}(y)}$
$=\mathbb{E}_{y\sim g}\log\frac{a(y)b(y)}{Z}\frac{Z^{\prime}}{a^{\prime}(y)b(y)}$
$=\mathbb{E}_{y\sim g}\log\frac{a(y)}{a^{\prime}(y)}-\log\frac{Z}{Z^{\prime}},$
where the expectations relative to $g$ can be estimated by obtaining samples from $g$ through sampling from $a$ and filtering with $b$, and where the values of $Z$ and $Z^{\prime}$ can be obtained by exploiting the equalities $Z=AR_{a},Z^{\prime}=AR_{a^{\prime}}$ (see App. B.5) and by estimating these acceptance rates.
In practice, these computations were performed using the disco [19] toolkit [^18] https://github.com/naver/disco and the estimations are made using $2^{20}$ (i.e., approximately 1M) samples from $a$. After filtering with $b$, we obtained around 2,300 samples from $g$ in the lexical constraints experiment ($AR_{a}=0.0023$) and 5,000 samples from $g$ in the sentiment reversal experiment ($AR_{a}=0.005$).
@@
Appendix E Lexical Constraint Experiment
@@
Comparison to a heuristic sampler
We performed a simple experiment using the heuristic sampler for keyword enforcement described in the Enforcement example paragraph of App. B.2, using “amazing” as the keyword to include in a 30-token generated text as in the experimental setting of Section 4.1. We define a sampler $s$ that checks whether we have already produced the string “amazing” within the first 29 tokens, and if not, forces the generation of “amazing” as the last token. With the Gemma-2B tokenizer, the string “amazing” can be generated with the single token [amazing] (along with [ama, zing], etc.), and we can easily design the sampler to produce this token as the 30th token if “amazing” does not appear earlier. For such a sampler $s$, we found that $\text{KL}(g||s)=6.06$, which is a significantly larger divergence than $\text{KL}(g||g^{\prime})=0.633$ for warm-start DPG-based $g^{\prime}$ as reported in Figure 4. More generally, the divergence $\text{KL}(g||s)$ is much greater than the divergences obtained for any $g^{\prime}$ based on DPG or SFT, whose values are always smaller than 3. This result appears to be consistent with the behavior we had anticipated in App. B.2, namely the poor ability of $s$ to approximate $g$ due to the biased positioning of “amazing” in the generated sequence.
@@
Appendix F Sentiment Reversal Experiment
@@
Appendix G Discussion of Additional Sampling Techniques: IMH and QRS
@@ G.2 Experimental Results
In the case of QRS and especially IMH, the quality gains come at a high cost in terms of additional inference time: to halve the projected KL divergence — which is a lower bound of the true KL divergence, see Footnote 24 — QRS requires 1.36 times the inference cost of Algorithm 3, while IMH requires 10 times the cost. Moreover, to reduce the KL divergence to 10% of the original KL, QRS uses almost 10 times the inference cost of Algorithm 3, while IMH consumes 50 times this cost. This experiment demonstrates that the random walk nature of MCMC methods can struggle to efficiently mimic the gold distribution. This result highlights why accurately approximating $g$ with the proposal distribution is crucial: it can lead to more efficient sampling that diverges less from the target distribution without incurring the extreme computational costs associated with methods like IMH.